OpenStudy (anonymous):

Solve the following equation algebra 1 -5=s-3

OpenStudy (anonymous):

Add 3 to each side to get s by itself first off.

OpenStudy (anonymous):

It would look like this: -5 = s-3 +3 +3 (-5+3)=-2 -2 = s

OpenStudy (anonymous):

Thanks